F(x) = 4x^3 - 5x 2x^(1/2). Determine the slope of the tangent when x = 4?
Answer by nyc_function(2741)

This is a calculus 1 question.

You need to find the derivative of F(x).

I'm looking for F'(x). So, I will do this by differenting termwise.

(4x^3)' = 12x^2

(-5x)' = -5

(2x^1/2)' = x^(-1/2)

We can write x^(-1/2) as 1/sqrt{x}.

So, F'(x) = 12x^2 + 1/sqrt{x} - 5

To find the slope of the tangent line, we now replace every x you see with 4 and simplify.

F'(4) = 12(4)^2 + 1/sqrt{4} - 5

F'(4) = 12*16 + 1/2 - 5

F'(4) = 375/2

The slope of the tangent line is 375/2.